• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name X. Qian

    Log entry time 07:58:37 on September24,2009

    Entry number 291367

    keyword=Owl shift summary

    00:00 keep taking low current data.

    RuN 12618

    00:20 am, asked MCC for 100 uA.

    Run 12619 Junk, just get 100 uA back,

    00:25 am turn on dithering start run 12620

    RUN 12620 00:28 ended 00:43

    1:00 turned off dithering.

    3:15am, beam is back as 98 uA.
    RUN 12624, data were taken during the tuning of beam position.

    RUN 12625 beam position is good now. 3:20 am, the beam position will need to stay at BPMA y 1.223 to avoid ion chamber trip.

    3:35 am, the beam problem came back again.

    5:15 am, MCC called, the highest beam is 75uA for now.
    we paged RC.

    RUN 12628, first good run in this shift @ 75uA finished @ 6:13 am
    RUN 12629, start 6:14 am. end 7:10am
    RUN 12630 start 7:10 am. end 7:40 am

    Turn off FBB, try to reach 100 uA.

    RUN 12631 start 7:40 am.